OpenStudy (calculusxy):

Emelia discovered that if she takes three-fourths of her age and adds 9, it produces the same result as when she takes one-fourth of her age and adds 21. How old is Emelia?

OpenStudy (calculusxy):

I have the equation for this as: \[\frac{ 3 }{ 4 }x+9=\frac{ 1 }{ 4 }x+21\]

OpenStudy (confluxepic):

Did you try subtracting 9 on both sides and going from there?

OpenStudy (calculusxy):

I did try but then I had 3/4x left. I didn't know how to isolate x with that.

OpenStudy (confluxepic):

Divide by 3/4 on both sides. Right?

OpenStudy (calculusxy):

Wait... I want to show you what I did so that you can catch the mistake. Is that okay?

OpenStudy (confluxepic):

Ok.

OpenStudy (calculusxy):